Primara Corporation has a standard cost system in which it applies overhead to products based on the standard direct labor-hours allowed for the actual output of the period. Data concerning the most recent year appear below:

Total budgeted fixed overhead cost for the year $ 515,900
Actual fixed overhead cost for the year $ 509,000
Budgeted direct labor-hours (denominator level of activity) 67,000
Actual direct labor-hours 68,000
Standard direct labor-hours allowed for the actual output 65,000

Required:

1. Compute the fixed portion of the predetermined overhead rate for the year. (Round Fixed portion of the predetermined overhead rate to 2 decimal places.)

2. Compute the fixed overhead budget variance and volume variance. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input all amounts as positive values.)
Answer #1

Fixed portion of predetermined Overhead rate

= Budgeted fixed oh/Budgeted direct labor hours

= 515,900/67,000

= 7.70

Fixed overhead budget variance

= Actual - Budgeted

= 509,000-515,900

= 6900 Favourable

​​​​​​Fixed overhead volume variance

= 515,900 - (65,000*7.70)

= 15,400 Unfavorable